OpenStudy (john_doe02125):

If h(x) = x − 1 and j(x) = −4x, solve h[j(5)]

OpenStudy (eazye_the_best):

If h(x) = x − 1 and j(x) = −4x, solve h[j(5)]

OpenStudy (eazye_the_best):

Now, we can find h[j(5)], which means "Plug x = j(5) into the h equation". Well, we know that j(5) is just -20. So we will plug in x = -20 into the h equation. This gives us: h[j(5)] = h[-20] = -20 - 1 = -21. This is our final answer.

OpenStudy (john_doe02125):

@EazyE_The_Best Thank you. I have one more question, if you could help me with it, it'd be great/

OpenStudy (eazye_the_best):

sure